PLEASE WRITE CLEAR AND READABLE ANSWERS. SHOW ALL WORK WHERE APPROPRIATE.

PLEASE WRITE CLEAR AND READABLE ANSWERS. SHOW ALL
Answer #1

a) V = ax + by^2

and E = dv/dr

Ex = d(ax + by^2)/dx = a

Ey = d(ax+ by^2)/dy = 2by

E = a i + 2by j N/C


b) at origin,

E = ai + 0j = a

F = qE = aq

c) dW = F.dr = qEdr

= q(a i + 2by j)dr

integrating,

W = qax + qby^2

x is from 0 to 0

y is from 0 to 2

W = 0 + qb(2^2 - 0 ) = 4qb